B. Evaluate the following algebraic expressions for the given value of x

1.2x—7 where x=4

Solution:

2.4x—8/3x+4 where x=—1

Solution:

3.x²—6x+5 where x=—3

Solution:


1
Expert's answer
2021-05-10T16:44:17-0400

1. 2x-7

"\Rightarrow \text{ At } x=4 \Rightarrow 2(4)-7=8-7=1"


"2. 4x-\dfrac{8}{3}x+4"

 at x=-1,

 "4(-1)-\dfrac{8}{3}(-1)+4=\dfrac{8}{3}"


3."x^2-6x+5"

  at x=-3, 

 "(-3)^2-6(-3)+5\n\n =9+18+5=32"